OpenStudy (calculusxy):

MEDAL!!!! As part of a fitness plan, Sabir stretches for 5 minutes before he walks on a treadmill at the same speed for 15 minutes every morning. The table below shows the calories he burns over time. Time in minutes .... Number of calories burned 5 - 25 10 - 45 15 - 65 20 - 85 Part A. Write an equation to represent the relationship between the time Sabir walks and the number of the calories he burns. Use x as the independent variable and y as the dependent variable. __________

OpenStudy (calculusxy):

@ganeshie8

OpenStudy (calculusxy):

@mathmale

OpenStudy (mathmale):

Convert the following data to points: 5 - 25 => (5,25) (the units of measurement are "minutes" and "calories burned" 15 - 65 => (15,65) We're going to find the equation of the straight line on which all four of the given points lie. First, please use the slope formula to find the slope of this line through the above 2 points. Now we're going to find the y-intercept, (0,b), of the line thru these four points: The general equation (slope-intercept form) is y=mx + b Substitute the value for the slope, m, that you've already found. Pick any one of the four given points, and substitute the x- and y-values of that point into y = m x + b. This will leave only the variable, b. Write the equation of this straight line using the slope you've already calculated and this b value. That's it.

OpenStudy (calculusxy):

@mathmale I am only in seventh grade and we have not yet talked about the the slope intercept formula yet. And neither do I think that this question requires slopes. I wrote the equation as to being: 4x + 5 = y

OpenStudy (calculusxy):

@hartnn

hartnn (hartnn):

ok, lets try it other way, since we know the equation would be linear, we could take the equation as y= mx+b since x=5, y =25 satisfies this equation, we get 25 = 5m +b got what i did ?

OpenStudy (calculusxy):

Is my equation correct or incorrect?

hartnn (hartnn):

and by the way, 4x+5 = y is absolutely correct :)

OpenStudy (calculusxy):

Thanks a lot :)

hartnn (hartnn):

welcome ^_^